OpenStudy (anonymous):

A triangle has one side that measures x, and the other two sides each measure 4 inches less than x. The perimeter is 19 inches. What is the measure of x? 2 3 5 9

OpenStudy (anonymous):

x+4-x+4-x=19

OpenStudy (anonymous):

19 = x + 2(x-4)

OpenStudy (anonymous):

expands out to 19 = x + 2x - 8 ----> 19 = 3x - 8

OpenStudy (anonymous):

x=9